OpenStudy (anonymous):

what are the points where 4x-y=9 and x-3y=16 intersect thanks

jhonyy9 (jhonyy9):

y=4x-9 x=16+3y --- this push in the first equation and will get y=4(16+3y)-9 y=64+12y-9 -11y=55 y=55/(-11) =-5 y=-5 x=16+3y x=16+3(-5) x=16-15 x=1 x=1 y=-5